OpenStudy (anonymous):

The absolute value of which of the following is smaller with E being very very small number greater than zero? a. [(4-E/2)^1/2]-2 b.[(4+E/2)^1/2]-2 I know the answer is b, but how do you figure it out without using a calculator?

OpenStudy (anonymous):

right, but your getting that the absolute values are the same. I need to know which absolute value is smaller

OpenStudy (phi):

re-reading it, The absolute value is the distance from 0, so we need 3 terms \[ \sqrt{1 \pm x} \approx 1 \pm \frac{x}{2} - \frac{x^2}{8} ....\] for 1+x we will get +x/2 - x^2/8 for 1-x -x/2 - x^2/8 the 2nd will have a slightly larger absolute value

OpenStudy (anonymous):

Got it, thank you :)